OpenStudy (anonymous):

Find the intervals of concavity and the inflection points of f(x)= x sqrt(x+3). I found my first derivative but I don't know if I'm doing the second derivative correctly. Help please!

OpenStudy (kainui):

Sure, so first off, do you understand why taking a derivative of anything would tell you where to find concavity and inflection points?

OpenStudy (kainui):

Just explain as best as you can what you think you're doing and I'll help you out! =D

OpenStudy (tkhunny):

Please SHOW your 1st Derivative.

OpenStudy (anonymous):

Here's my first derivative \[\frac{ 3(x+2)}{ 2\sqrt{x+3}}\]

OpenStudy (anonymous):

okay when I solve to get my second derivative I get this. I don't know if it's right so far. \[\frac{ 12x+36 }{ \sqrt{x+3}} - \frac{ 6x+2 }{ \sqrt{x+3}} = \frac{ 6x+38 }{\sqrt{x+3} }\]

OpenStudy (tkhunny):

Not really sure how you managed that 2nd Derivative. Can you describe your methodology? Quotient Rule? What?

OpenStudy (anonymous):

I used the quotient rule. Oh but I just realized I forgot the \[g^2\]from the quotient rule so my answer must be wrong

OpenStudy (tkhunny):

Should be: \(\dfrac{3}{2}\dfrac{\sqrt{x+3}\cdot\dfrac{d}{dx}(x+2) - (x+2)\cdot\dfrac{d}{dx}\sqrt{x+3}}{x+3}\). Please note how removing the constants up front made life a whole lot easier.

OpenStudy (anonymous):

Thats makes things a lot clearer! But how did you remove the radical from \[\sqrt{x+3}\] in the bottom.

OpenStudy (tkhunny):

\(g^{2}\), remember?

OpenStudy (anonymous):

Oh right! okay, let me solve it now to see if im doing it right.

OpenStudy (anonymous):

\[\frac{ 3 }{ 2}\frac{ \sqrt{x+3}-(x+2)(\frac{ 1 }{ 2 }) (x+3)^{-\frac{ 1 }{ 2 }}}{{x+3} }\] Okay so i know in order to make \[(x+3)^{\frac{ -1 }{ 2}}\] I would have to bring it to the bottom would that be \[x+3\sqrt{x+3}\] ?

OpenStudy (anonymous):

I don't know what to do after this.

OpenStudy (tkhunny):

Good try. You can work on your bookkeeping, later. You should get: \(\dfrac{3}{4}\cdot\dfrac{x+4}{(x+3)^{3/2}}\) for the 2nd Derivative.

OpenStudy (anonymous):

Hmm, okay so from there how would I find the inflection point?

OpenStudy (tkhunny):

What makes the 1st Derivative 0 or undefined? What makes the 2nd Derivative 0 or undefined?

OpenStudy (anonymous):

for the first derivative I got x= -2 is that correct?

OpenStudy (anonymous):

and the second derivative x=-4?

OpenStudy (tkhunny):

We should point out that the Domain of f(x) is \(x \ge -3\). Do you see why this is so? How does this make you feel about x = -4 for a zero of the 2nd derivative?

OpenStudy (anonymous):

Why is the domain of\[f(x) \ge -3\] ?

OpenStudy (tkhunny):

\(\sqrt{x+3}\) Does x = -5 work in there?

OpenStudy (anonymous):

No it would make it undefined.

OpenStudy (tkhunny):

So, what is the Domain? What values work in that square root?

OpenStudy (anonymous):

-2 would work?

OpenStudy (anonymous):

or does it have to equal to zero?

OpenStudy (tkhunny):

It's continuous. \(x \ge -3\)

OpenStudy (anonymous):

OH! okay I understand.

OpenStudy (anonymous):

so then -4 would not work for the bottom. but when i set the top to 0 it gives me x= -4

OpenStudy (tkhunny):

Yup. It's not in the Domain. Just ignore it. Next important question: What is the Domain of \(f'(x)\)?

OpenStudy (anonymous):

Would it be the same thing ? \[x \ge -3\]

OpenStudy (tkhunny):

Put x = -3 in there and see how that goes.

OpenStudy (anonymous):

It gives me 0

OpenStudy (tkhunny):

For the DENOMINATOR! Is that okay?

OpenStudy (anonymous):

hmm so then is it just \[x >-3\] ?

OpenStudy (tkhunny):

That is correct. Same thing for the 2nd Derivative. Recap: f(x) Domain \(x \ge -3\) f'(x) Zero: x = -2 Fails to exist: x = -3 f"(x) Zero: None Fails to exist: x = -3 We are ready to identify the Critical Points. What are they?

OpenStudy (anonymous):

x = -2 ?

OpenStudy (tkhunny):

And x = -3. Don't miss these oddballs.

OpenStudy (anonymous):

Oh okay. Got it.

OpenStudy (tkhunny):

Okay, one last thing. Classify the Critical Points.

OpenStudy (anonymous):

Do I plug them into my second derivative?

OpenStudy (anonymous):

Or my original function?

OpenStudy (tkhunny):

No, that is not likely to work, since x = -3 won't go! Think about them. x = -2 f'(-2) = 0 -- This this could be a min or a max. f"(-2) > 0 This is a minimum value. It turns out to be a Global minimum. The second derivative is everywhere positive, so this graph won't be doing any more tricks. x = -3 What do you think?

OpenStudy (anonymous):

So x= -3 can't be a minimum or a maximum?

OpenStudy (tkhunny):

Not so. The derivative doesn't tell you anything, but it still can be. You must check places like this by yourself. What is the sign of the derivative just shy of x = -3? We already established that the 2nd derivative is positive everywhere it exists.

OpenStudy (tkhunny):

We often encounter this kind of problem with artificial limits on the Domain. The problem statement suggests a function, but then allows x only between two values, say 1 and 2. The derivative knows nothing of the artificial limitations. You must examine these oddities by yourself. This problem has a more natural endpoint, but the havoc with the derivatives is the same.

OpenStudy (tkhunny):

f(-3) = 0 f'(-3+a hair) < 0 f"(-3+a hair) > 0 This guy is a local maximum.

OpenStudy (anonymous):

Now I understand. So would it be concave up on \[(-3, \infty)\] ?

OpenStudy (tkhunny):

Yes.

OpenStudy (anonymous):

Wow you helped me so much with this. I actually understand it now. I just have to work on finding the second derivative. Thank you so much for your help and time. I really appreciate it! :)

OpenStudy (tkhunny):

No worries. It's what we strive to do around here. Good working hanging in there!